An object is thrown from from the ground upward with an initial speed of 3.75 m/s. How long will the object be in the air before it lands on the ground?



1
Expert's answer
2021-07-25T09:15:55-0400

Let's first find the time that the object takes to reach its maximum height:


"v= v_0-gt,""0= v_0-gt,""t=\dfrac{v_0}{g}=\dfrac{3.75\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=0.38\ s."

Finally, we can find the flight time of the object:


"t_{flight} = 2t=2\cdot0.38\ s=0.76\ s."